Dashboard

All systems are operational.

Boxed IT Status

Overview of the operational status of the Boxed IT server and network infrastructure. All major incidents will be reported on this page and on twitter. Please follow @boxedit on twitter for the latest updates.

Planned Maintenance

  • Mirror server maintenance

    5 years ago —
    In Progress

    The mirror server will be down for a short period due to system reorganisation and operating system upgrades.

  • Virtual Machine Platform maintenance at Brussels 1 Facility

    5 years ago —
    In Progress

    Due to server hardware maintenance, the following services will be interrupted:

    • Hosted websites on ares.boxed-it.com
    • Hosted virtual machines at Brussels 1 facility
    • Incoming mail, outgoing mail, webmail

    DNS Name server service will remain operational, and no incoming e-mail will be lost - it will be queued for delivery.

    Update 22:15: Maintenance commencing

    Update 23:40: Maintenance completed, all customer virtual machines are starting up. Mail, DNS and web services are online.

  • VPS platform maintenance

    5 years ago —
    In Progress

    We will be performing hypervisor maintenance on the physical host machine serving your Virtual Private Server (VPS) to better serve you in the future.

    Impact

    Your VPS will be shut down when maintenance starts, and will be started again after maintenance completes.

    Maintenance window

    • Start: 02 January 2020 - 23:00 CET
    • End: 03 January 2020 - 02:00 CET

    Please note it is likely the maintenance will not take the entire window.

Past Incidents

No incidents reported.

No incidents reported.

No incidents reported.

No incidents reported.

No incidents reported.

No incidents reported.

9 years ago —
Fixed
9 years ago —

Our primary router crashed again, as on July 10th, causing slow convergence and recovery for the rest of the network.

Update 19:15: The router has been shut down and the network has been stabilized. Our team is on the way to the datacenter to replace the router with a spare model.

Update 20:45: The router power supply has been replaced and the configuration has been adjusted to demote the router to secondary until we are convinced that the hardware issue has been resolved.

Update 22:08: Configuration has been restored, close monitoring of the router will continue.